OpenStudy (anonymous):

Solve: x − 4 = −15 I did this equation and I got the answer to be 19 would this be correct

Parth (parthkohli):

Did you skip through the negative sign right there?

OpenStudy (anonymous):

Check it once more..

Parth (parthkohli):

\( \color{Black}{\Rightarrow x - 4 = \mathbf{-}15}\)

OpenStudy (anonymous):

x − 4 = −15 x=-15+4=-11

Parth (parthkohli):

\( \color{Black}{\Rightarrow x = -15 + 4}\)

Parth (parthkohli):

@sai1234 Please don't post answers wherever possible. Thank you.

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

X -11

Parth (parthkohli):

Yes.